OpenStudy (anonymous):

Which is an equation for the line with slope 1/3 through (-2, 1)?

OpenStudy (anonymous):

Given (x1, y1) = (-2, 1) m= 1/3 Eq of line in slope -point form is given as: y-y1 =m(x-x1) y-1=1/3[x-(-2)] y-1 =1/3 (x+2) 3y-3= x+2 3y= x+2+3 3y= x+5 is the required equation. @savannah123456789